# Bright traveling breathers in media with long-range, nonconvex dispersion

### v2

@@ III.2 Parametrization and numerical continuation
Having described the computation of a numerically accurate waveform in the weakly nonlinear regime, we now lay out the continuation procedure to obtain a family of solutions from this known solution. Implicit to the product basis representation in Eq. (15) for spatially localized unit mean breathers, we have a two-parameter characterization (envelope velocity $c$ and time-period $T$). Given a known solution, we perform a line search for a fixed time-period $T$, by varying the velocity ($c$) parameter (referred to as $c$-continuation). We perform several such line searches starting from different weakly nonlinear solutions. The search algorithm at the $j^{th}$ $c$-continuation step is
${\textbf{P}}_{1n}^{\dagger(j)}{\textbf{P}}_{1n}^{(j)}\Delta\beta$ $=-{\textbf{P}}_{1n}^{\dagger(j)}{\textbf{P}}_{0}^{(j)}\beta_{n};n=1,2,\cdots$ (19)
where the iterations are seeded with a known solution ($\beta^{(j-1)}$, $c^{(j-1)}$), while a traveling breather with fixed velocity $c^{(j)}=c^{(j-1)}+\delta c$ is sought. Note that similar space-time boundary value computations have been used to compute breathers and traveling breathers in discrete lattice systems James (2018) and modified NLS-type models Ward et al. (2019, 2020).
@@ IV.1 BBM bright traveling breather solutions
We compute five branches of BBM traveling breathers, bifurcating from the NLS limit. The edge of the focusing regime ($\sigma>0$ in Eq. (11)) of the NLS reduction is marked by the inflection point of the unit mean linear dispersion relation (3) $\tilde{k}=\sqrt{3}\approx 1.73$. It is crucial to initialize the Newton iterations with an accurate initial guess in the nearly monochromatic regime for convergence to traveling breathers. In order to use a NLS bright soliton as the initial seed for the continuation procedure, it is necessary to initialize the carrier wavenumber $\tilde{k}>\sqrt{3}$. The five traveling breather branches are characterized by the carrier wavenumbers
@@
We refer to each of these branches as Family-1, 2, 3, 4 and 5, respectively. We perform the computations on large domains for each of the breather families 1-5, with $2L\in\{400,200,500,500,500\}$ and respectively. The spatial discretization for all the computed families is $\Delta x={2L}/{2M}=0.08$, while $N=16$ (32 Fourier modes) is found to be an appropriate discretization in time.The NLS approximations provide good initial guesses for computing traveling breathers with carrier wave numbers $\tilde{k}>2.4$.
The initial guess for NCG iterations is the bright soliton solution of the NLS equation Eq. (11), where $\gamma(\tilde{k})=(5\tilde{k}^{2}+3)/(6\tilde{k}^{3}+18\tilde{k})$ and $\omega_{0}$ (Eq. (3)), $\partial_{\tilde{k}}\omega_{0}$, $\partial_{\tilde{k}\tilde{k}}\omega_{0}$ are evaluated at $\tilde{k}=\tilde{k}^{(j)}$ for some $j$ and $u_{0}=1$. We initialize the computations with $\tilde{a}\approx 0.15$. The NLS bright soliton envelope is localized, whereas the computed traveling breathers are found to exhibit oscillatory tails, albeit very small for small $\tilde{a}$. It is convenient to introduce the traveling breather frequency in the co-moving frame $\Omega=2\pi/T$.
Figure 1: (a) A traveling breather in the weakly nonlinear regime (solid blue line) of Family-1 with $T\approx 18.180$, $c\approx-0.031$ on a spatial domain with $L=400$. The numerical envelope (dashed black line) is compared to the NLS bright soliton envelope (solid red line). (b) Evaluation of the traveling breather at $\chi=0$ (in blue) compared with the leading order NLS approximation (red solid line, Eq. (16)).
In Fig. 1, we compare the computed and NLS bright soliton profiles for the first traveling breather of Family-1. The profiles agree to within $1\%$. The numerical envelope profiles are extracted using the Hilbert transform Nakayama and Lamb (2020). Furthermore, the traveling breather frequency $\Omega$ and velocity $c$ are within $1\%$ of the bright soliton predictions (16). The first computed solution for all traveling breather families exhibits similar agreement to the NLS bright soliton prediction (16).
To recover other traveling breathers, we apply $c$-continuation. The continuation procedure is found to introduce a shift in the wave mean. At the end of $c$-continuation for each family, we scale all the obtained traveling breathers to unit mean by utilizing Eq. (2). Consequently, although we fix the time period $T$ during continuation, the rescaling to unit mean implies that the time period of the computed traveling breathers changes across each family.
Performing $c$-continuation, we observe an increase in traveling breather amplitude with decreasing $c$. Define the traveling breather nonlinear frequency shift
@@
where $\Omega_{0}=\omega_{0}-\tilde{k}\partial_{\tilde{k}}\omega_{0}$ is the linear frequency in the co-moving frame. In Fig. 2(a), the prediction $\Delta\Omega=\epsilon^{2}/2$ from Eq. (16) is compared with the computed frequency shift (21) for (Family-1) across a range of amplitudes. As expected, there is good agreement at low amplitudes and deviation at large amplitudes.
Figure 2: Traveling breather frequency shift from computation (dots) and predicted (solid) traveling breathers.
Figure 3 displays six traveling breather solutions from Family-1 at $\tau=0$. To the eye, the solution in Fig. 3(a) appears localized, decaying to 1 as $|\chi|\to\infty$. In fact, all computed traveling breathers exhibit an oscillatory background. This is consistent with rigorous studies of breathers in Klein-Gordon equations where it was proven that, in the non-integrable case, small amplitude breather solutions are accompanied by exponentially small oscillatory tails Segur and Kruskal (1987); Soffer and Weinstein (1999). As the traveling breather amplitude increases in Fig. 3, the oscillatory background becomes more prominent, eventually reaching an amplitude that is comparable to the traveling breather itself. The increase in carrier amplitude is accompanied by a narrowing of the traveling breather’s width such that, in Figs. 3(e,f), the traveling breather itself is narrower than the carrier wavelength.
Figure 3: Family-1 of traveling breather solutions of the BBM equation with $(T,c)\approx$ (a) $16.917,-0.036$, (b) $11.969,-0.076$, (c) $11.942,-0.079$, (d) $11.906,-0.082$, (e) $11.316,-0.099$, (f) $10.682,-0.129$.
We point to some other general trends within the computed solution families. The upper limit to the traveling breather velocity $c$ is approached in the vanishing amplitude regime so that it is the linear group velocity for Family-$j$ $c<\partial_{\tilde{k}}\omega_{0}(\tilde{k}^{(j)},1)$, implying that all traveling breather velocities are negative. On the other hand, the computations do not apparently indicate a lower bound to the breather velocities:
@@
This implies that, when bifurcating from the NLS bright soliton, traveling breather velocities are always negative $c\in(-\infty,0)$.
Defining the traveling breather envelope amplitude as
$\tilde{a}=\max_{\tau\in[0,T]}\beta(0,\tau)-\min_{\tau\in[0,T]}\beta(0,\tau),$ (23)
we observe that the traveling breather width narrows relative to the carrier wavelength. These trends are shown in Fig. 3.
@@
and the carrier wavenumber $k$ (cf. Eq. (6)), with the carrier frequency $\omega=\omega(k,a)$ uniquely determined by the unit mean ($\overline{u}=1$) constraint (7). Since the traveling breather frequency in the co-moving frame is known $\Omega=2\pi/T$, we therefore have the additional compatibility relation
@@
Traveling breathers impart a phase shift to the periodic background, which we normalize to $\sigma\in[-\pi,\pi]$. The phase shift quantifies the amount by which the periodic background has advanced/receded across the traveling breather core. In our computations, the phase shift is implicitly determined by the spatial domain $[-L,L]$ and the imposition of spatially periodic boundary conditions. The number of carrier wavelengths that fit in the domain is $N=\lfloor Lk/\pi\rfloor$ so that the difference $\Delta\chi=2L-2\pi N/k\geq 0$ represents the spatial mismatch. The normalized phase shift is determined according to
@@
For our traveling breather computations, we input the four parameters $(c,T,L,\overline{u}=1)$. The two relations (25) and (26) can be used to determine $a=a(k)$ and $\sigma=\sigma(k)$. Thus, unit mean traveling breathers constitute a three-parameter family of solutions.
@@
Figure 4: Variation of the (a) periodic background amplitude (23), (b) breather amplitude (24), (c) periodic background wavenumber, (d) phase shift $\sigma$ (26) and (e) time-period $T$ with breather velocity $c$ for Family-1 of BBM breathers.
@@
In Fig. 4 we show how the parameters $a$, $\tilde{a}$, $k$ , $\sigma$ and $T$ vary with the breather velocity $c$ for Family-1. As $c$ is decreased, the parameters $a$, $\tilde{a}$ and $k$ increase, with $k$ (and as a consequence $a(k)$) rapidly approaching an asymptotic value in the strongly nonlinear regime. The phase shift $\sigma(k)$ also limits to an asymptotic value for large amplitude (more negative $c$) traveling breathers, while displaying both positive and negative phase shifts across the range of velocities. By continuity, there is a traveling breather exhibiting a zero phase shift. Finally, we remark that, since we initialized the computations of Family-1 with the NLS bright soliton (16) where $\tilde{a}\approx 0.15$, the largest value of $c$, $c_{\rm max}\approx-0.031$, is less than the theoretical upper bound (22) $\partial_{\tilde{k}}\omega_{0}(\tilde{k}^{({1})},1)\approx-0.0308$, which applies in the limit $\tilde{a}\to 0$. We did not attempt to compute smaller amplitude solutions. The smallest velocity for which we compute a traveling breather solution is $c\approx-0.129$ and is shown in Fig. 3(f). At this point, the existence of a lower bound for breather velocities is unknown but remains interesting for future investigations.
Figure 5: Family-5 of BBM traveling breathers. (a) $(\tilde{a},c,T)\approx 0.162,-0.110,10.142$, (b) $(\tilde{a},c,T)\approx 0.306,-0.124,9.708$, (c) $(\tilde{a},c,T)\approx 0.623,-0.133,9.486$.
@@
Figure 6: (a) Projections in amplitude of the velocity ($c$)-frequency ($\Omega$) relation of the five traveling breather branches. (b) Computed unit mean traveling breathers from the five families (filled circles) and the weakly nonlinear prediction $T=\mathcal{T}$ (28) (colored surface).
@@
The fifth branch of computed traveling breathers crossing the linear inflection point shown in Fig. 5 (a)-(c) merits special mention. Given their proximity to the zero dispersion line ($\Omega_{0}=3\sqrt{3}/8$) in the weakly nonlinear regime, they display a prominent periodic background even for small amplitudes. For amplitudes $0.16\leq\tilde{a}\leq 0.62$, the carrier wavenumber of the breather core is found to lie in the interval $\tilde{k}\in[1.67,2.42]$, which is below the linear dispersion inflection point $\tilde{k}=\sqrt{3}\approx 1.73$. When $\tilde{k}<\sqrt{3}$, the NLS equation (11) is repulsive/defocusing. This persistence of bright traveling breathers across the zero-dispersion line into the defocusing regime is an intriguing feature, suggesting the need for a higher order NLS model to describe them Kivshar and Afanasjev (1991); Kivshar and Luther-Davies (1998).
@@
In order to assess how close to the NLS regime computed traveling breathers are, we plot the relationship between the traveling breather frequency in the co-moving frame with frequency $\Omega$ and velocity $c$ for each solution family in Fig. 6(a). To compare with the NLS bright soliton (16), we obtain a relationship between the linear frequency in the co-moving frame $\Omega_{0}=2\tilde{k}^{3}/(1+\tilde{k}^{2})^{2}$ and the linear group velocity $c_{0}=\partial_{\tilde{k}}\omega_{0}=(1-\tilde{k}^{2})/(1+\tilde{k}^{2})^{2}$ by eliminating $\tilde{k}$ to obtain
$\Omega_{0}(c_{0})={2\sqrt{-2c_{0}}\frac{((2c_{0}+1)+\sqrt{8c_{0}+1})^{3/2}}{(1+\sqrt{8c_{0}+1})^{2}}},$ (27)
@@
where $-1/8<c_{0}<0$ in the focusing regime. The computed traveling breathers lie close to the $\Omega_{0}(c_{0})$ curve in Fig. 6(a) for small amplitudes, and depart from the curve for larger amplitudes or fewer carrier wave oscillations in the breather core. We conclude this section with a concise representation of all the computed traveling breather families in Fig. 6 (b). To this end, a convenient set of defining parameters is the traveling breather amplitude $\tilde{a}$, the velocity $c$, and the time period in the co-moving frame $T$ (or the associated angular frequency $\Omega=(2\pi)/T$). From Eqs. (16) and (21) with $\Delta\Omega=\epsilon^{2}/2$, the time-period $T$ in the NLS approximation is
@@
where $\gamma(c)=\frac{\sqrt{-2c}}{6\sqrt{(2c+1)+\sqrt{8c+1}}}\frac{-5-4c-5\sqrt{8c+1}}{4c-1-\sqrt{8c+1}}$ and $\tilde{k}(c)\equiv\sqrt{\frac{-(2c+1)-\sqrt{8c+1}}{2c}}$ is determined by inverting $c=\partial_{\tilde{k}}\omega_{0}(\tilde{k},1)$. All NLS-like envelope solitons, enclosing several carrier wave oscillations within the breather core lie close to the two-parameter surface $T=\mathcal{T}(\tilde{a},c)$. These breather cores decay rapidly to a very small amplitude ($a\ll 1$) periodic background. This surface is depicted in Fig. 6 (b). The parameters $(\tilde{a},c,T)$ associated with each computed traveling breather are also rendered in the figure. In the weakly nonlinear regime, the traveling breathers of all five families reside close to $T=\mathcal{T}(\tilde{a},c)$. Eventually, they depart from the surface, exhibiting a larger time period than weakly nonlinear theory predicts. For the families $1$-$4$, the strongly nonlinear breathers limit to enclosing very few carrier oscillations within the breather cores, with the envelope widths being comparable to the cnoidal (background) carrier wavelength. An exception in this regard is the family-5; wherein the breather core possesses a slowly varying envelope despite a large amplitude $\tilde{a}$. As a final remark, we draw attention to families 2 and 3, computed on domains with $L=100$ and $250$ respectively and seeded with an identical NLS envelope soliton (see Figs.  6). The role of the computational domain length as an additional parameter (besides $c$ and $T$) is clear, as the scatter plots diverge noticeably beyond the weakly nonlinear regime.

### v3

@@ III.2 Parametrization and numerical continuation
Having described the computation of a numerically accurate waveform in the weakly nonlinear regime, we now lay out the continuation procedure to obtain a family of solutions from this known solution. Implicit to the product basis representation in Eq. (15) for spatially localized unit-mean breathers, we have a two-parameter characterization (envelope velocity $c$ and timeperiod $T$). Given a known solution, we perform a line search for a fixed timeperiod $T$, by varying the velocity $c$ parameter (referred to as $c$ continuation). We perform several such line searches starting from different weakly nonlinear solutions. The search algorithm at the $j$th $c$-continuation step is
${\textbf{P}}_{1n}^{\dagger(j)}{\textbf{P}}_{1n}^{(j)}\Delta\beta$ $=-{\textbf{P}}_{1n}^{\dagger(j)}{\textbf{P}}_{0}^{(j)}\beta_{n};n=1,2,_{\cdots},$ (19)
where the iterations are seeded with a known solution ($\beta^{(j-1)}$ and $c^{(j-1)}$), while a traveling breather with fixed velocity $c^{(j)}=c^{(j-1)}+\delta c$ is sought. Note that similar space-time boundary-value computations have been used to compute breathers and traveling breathers in discrete lattice systems James (2018) and modified NLS-type models Ward et al. (2019, 2020).
@@ IV.1 BBM bright traveling breather solutions
We compute five branches of BBM traveling breathers, bifurcating from the focusing NLS limit. The edge of the focusing regime [$\sigma>0$ in Eq. (11)] of the NLS reduction is marked by the inflection point of the unit-mean linear dispersion relation (3) $\tilde{k}=\sqrt{3}\approx 1.73$. It is crucial to initialize the Newton iterations with an accurate initial guess in the nearly monochromatic regime for convergence to traveling breathers. In order to use a NLS bright soliton as the initial seed for the continuation procedure, it is necessary to initialize the carrier wavenumber $\tilde{k}>\sqrt{3}$. The five traveling breather branches are characterized by the carrier wave numbers
@@
We refer to each of these branches as families 1, 2, 3, 4, and 5, respectively. We perform the computations on large domains for each of the breather families 1-5, with $2L\in\{400,200,500,500,500\}$, respectively. The spatial discretization for all the computed families is $\Delta x={2L}/{2M}=0.08$, while $N=16$ (32 Fourier modes) is found to be an appropriate discretization in time. The NLS approximations provide good initial guesses for computing traveling breathers with carrier wave numbers $\tilde{k}>2.4$.
The initial guess for NCG iterations is the bright soliton solution of the NLS equation (11), where $\gamma(\tilde{k})=(5\tilde{k}^{2}+3)/(6\tilde{k}^{3}+18\tilde{k})$ and $\omega_{0}$ [Eq. (3)], $\partial_{\tilde{k}}\omega_{0}$ and $\partial_{\tilde{k}\tilde{k}}\omega_{0}$ are evaluated at $\tilde{k}=\tilde{k}^{(j)}$ for some $j$ and $u_{0}=1$. We initialize the computations with $\tilde{a}\approx 0.15$. The NLS bright soliton envelope is localized, whereas the computed traveling breathers are found to exhibit oscillatory tails, albeit very small for small $\tilde{a}$. It is convenient to introduce the traveling breather frequency in the comoving frame $\Omega=2\pi/T$.
Figure 1: (a) Traveling breather in the weakly nonlinear regime (blue solid line) of family 1 with $T\approx 18.180$ and  $c\approx-0.031$ on a spatial domain with $L=400$. The numerical envelope (black dashed line) is compared to the NLS bright soliton envelope (red solid line). (b) Evaluation of the traveling breather at $\chi=0$ (blue solid line) compared with the leading-order NLS approximation  (16) (red solid line).
In Fig. 1 we compare the computed and NLS bright soliton profiles for the first traveling breather of family 1. The profiles agree to within $1\%$. The numerical envelope profiles are extracted using the Hilbert transform Nakayama and Lamb (2020). Furthermore, the traveling breather frequency $\Omega$ and velocity $c$ are within $1\%$ of the bright soliton predictions (16). The first computed solution for all traveling breather families exhibits similar agreement to the NLS bright soliton prediction (16).
To recover other traveling breathers, we apply $c$ continuation. The continuation procedure is found to introduce a shift in the wave mean. At the end of $c$ continuation for each family, we scale all the obtained traveling breathers to unit mean by utilizing Eq. (2). Consequently, although we fix the time period $T$ during continuation, the rescaling to unit mean implies that the time period of the computed traveling breathers changes across each family.
Performing $c$ continuation, we observe an increase in traveling breather amplitude with decreasing $c$. Define the traveling breather nonlinear frequency shift
@@
where $\Omega_{0}=\omega_{0}-\tilde{k}\partial_{\tilde{k}}\omega_{0}$ is the linear frequency in the co-moving frame. In Fig. 2(a) the prediction $\Delta\Omega=\epsilon^{2}/2$ from Eq. (16) is compared with the computed frequency shift (21) for family 1 across a range of amplitudes. As expected, there is good agreement at low amplitudes and deviation at large amplitudes.
Figure 2: Traveling breather frequency shift from computation (black dotted line) and predicted (red solid line) traveling breathers.
Figure 3 displays six traveling breather solutions from family 1 at $\tau=0$. To the eye, the solution in Fig. 3(a) appears localized, decaying to 1 as $|\chi|\to\infty$. In fact, all computed traveling breathers exhibit an oscillatory background. This is consistent with rigorous studies of breathers in Klein-Gordon equations where it was proven that, in the nonintegrable case, small-amplitude breather solutions are accompanied by exponentially small oscillatory tails Segur and Kruskal (1987); Soffer and Weinstein (1999). As the traveling breather amplitude increases in Fig. 3, the oscillatory background becomes more prominent, eventually reaching an amplitude that is comparable to the traveling breather itself. The increase in carrier amplitude is accompanied by a narrowing of the traveling breather’s width such that, in Figs. 3(e,f), the traveling breather itself is narrower than the carrier wavelength.
Figure 3: Family 1 of traveling breather solutions of the BBM equation with (a) $(T,c)\approx 16.917,-0.036$, (b) $(T,c)\approx 11.969,-0.076$, (c) $(T,c)\approx 11.942,-0.079$, (d) $(T,c)\approx 11.906,-0.082$, (e)$(T,c)\approx 11.316,-0.099$, (f) $(T,c)\approx 10.682,-0.129$.
We point to some other general trends within the computed solution families. The upper limit to the traveling breather velocity $c$ is approached in the vanishing amplitude regime so that it is the linear group velocity for family $c<\partial_{\tilde{k}}\omega_{0}(\tilde{k}^{(j)},1)$, implying that all traveling breather velocities are negative. On the other hand, the computations do not apparently indicate a lower bound to the breather velocities:
@@
This implies that, when bifurcating from the NLS bright soliton, traveling breather velocities are always negative, i.e. $c\in(-\infty,0)$.
The traveling breather envelope amplitude is defined as
$\tilde{a}=\max_{\tau\in[0,T]}\beta(0,\tau)-\min_{\tau\in[0,T]}\beta(0,\tau).$ (23)
In Fig. 3 we observe that the traveling breather width narrows relative to the carrier wavelength.
@@
and the carrier wavenumber $k$ [cf. Eq. (6)], with the carrier frequency $\omega=\omega(k,a)$ uniquely determined by the unit-mean ($\overline{u}=1$) constraint (7). Since the traveling breather frequency in the comoving frame is known $\Omega=2\pi/T$, we have the additional compatibility relation
@@
Traveling breathers impart a phase shift to the periodic background, which we normalize to $\sigma\in[-\pi,\pi]$. The phase shift quantifies the amount by which the periodic background has advanced or receded across the traveling breather core. In our computations, the phase shift is implicitly determined by the spatial domain $[-L,L]$ and the imposition of spatially periodic boundary conditions. The number of carrier wavelengths that fit in the domain is $N=\lfloor Lk/\pi\rfloor$ so the difference $\Delta\chi=2L-2\pi N/k\geq 0$ represents the spatial mismatch. The normalized phase shift is determined according to
@@
For our traveling breather computations, we input the four parameters $c,T,L,\overline{u}=1$. The two relations (25) and (26) can be used to determine $a$ and $\sigma$. Thus, unit-mean traveling breathers constitute a three-parameter family of solutions.
@@
Figure 4: Variation of the (a) periodic background amplitude (23), (b) breather amplitude (24), (c) periodic background wave number, (d) phase shift $\sigma$ (26), and (e) timeperiod $T$ with breather velocity $c$ for family 1 of BBM breathers.
@@
In Fig. 4 we show how the parameters $a$, $\tilde{a}$, $k$ , $\sigma$, and $T$ vary with the breather velocity $c$ for family 1. As $c$ is decreased, the parameters $a$, $\tilde{a}$, and $k$ increase, with $k$ and $a$ rapidly approaching an asymptotic value in the strongly nonlinear regime. The phase shift $\sigma$ also limits to an asymptotic value for large-amplitude (more negative $c$) traveling breathers, while displaying both positive and negative phase shifts across the range of velocities. By continuity, there is a traveling breather exhibiting a zero phase shift. Finally, we remark that, since we initialized the computations of family 1 with the NLS bright soliton (16) where $\tilde{a}\approx 0.15$, the largest value of $c$, $c_{\rm max}\approx-0.031$, is less than the theoretical upper bound (22) $\partial_{\tilde{k}}\omega_{0}(\tilde{k}^{({1})},1)\approx-0.0308$, which applies in the limit $\tilde{a}\to 0$. We did not attempt to compute smaller-amplitude solutions. The smallest velocity for which we compute a traveling breather solution is $c\approx-0.129$, which is shown in Fig. 3(f). At this point, the existence of a lower bound for breather velocities is unknown but remains interesting for future investigations.
Figure 5: Family 5 of BBM traveling breathers with (a) $(\tilde{a},c,T)\approx 0.162,-0.110,10.142$, (b) $(\tilde{a},c,T)\approx 0.306,-0.124,9.708$, and (c) $(\tilde{a},c,T)\approx 0.623,-0.133,9.486$.
@@
Figure 6: (a) Projections in amplitude of the velocity ($c$)-frequency ($\Omega$) relation of the five traveling breather branches. (b) Computed unit-mean traveling breathers from the five families (closed circles) and the weakly nonlinear prediction $T=\mathcal{T}$ (28) (colored surface).
@@
The fifth branch of computed traveling breathers crossing the linear inflection point shown in Figs. 5 (a)-5(c) merits special mention. Given their proximity to the zero-dispersion line ($\Omega_{0}=3\sqrt{3}/8$) in the weakly nonlinear regime, they display a prominent periodic background even for small amplitudes. For amplitudes $0.16\leq\tilde{a}\leq 0.62$, the carrier wavenumber of the breather core is found to lie in the interval $\tilde{k}\in[1.67,2.42]$, which is below the linear dispersion inflection point $\tilde{k}=\sqrt{3}\approx 1.73$. When $\tilde{k}<\sqrt{3}$, the NLS equation (11) is repulsive or defocusing. This persistence of bright traveling breathers across the zero-dispersion line into the defocusing regime is an intriguing feature, suggesting the need for a higher-order NLS model to describe them Kivshar and Afanasjev (1991); Kivshar and Luther-Davies (1998).
@@
In order to assess how close to the NLS regime computed traveling breathers are, we plot the relationship between the traveling breather frequency in the comoving frame with frequency $\Omega$ and velocity $c$ for each solution family in Fig. 6(a). To compare with the NLS bright soliton (16), we obtain a relationship between the linear frequency in the co-moving frame $\Omega_{0}=2\tilde{k}^{3}/(1+\tilde{k}^{2})^{2}$ and the linear group velocity $c_{0}=\partial_{\tilde{k}}\omega_{0}=(1-\tilde{k}^{2})/(1+\tilde{k}^{2})^{2}$ by eliminating $\tilde{k}$ to obtain
$\Omega_{0}(c_{0})={2\sqrt{-2c_{0}}\frac{[(2c_{0}+1)+\sqrt{8c_{0}+1}]^{3/2}}{(1+\sqrt{8c_{0}+1})^{2}}},$ (27)
@@
where $-\frac{1}{8}<c_{0}<0$ in the focusing regime. The computed traveling breathers lie close to the $\Omega_{0}(c_{0})$ curve in Fig. 6(a) for small amplitudes and depart from the curve for larger amplitudes or fewer carrier wave oscillations in the breather core. We conclude this section with a concise representation of all the computed traveling breather families in Fig. 6 (b). To this end, a convenient set of defining parameters is the traveling breather amplitude $\tilde{a}$, the velocity $c$, and the time period in the comoving frame $T$ (or the associated angular frequency $\Omega=2\pi/T$). From Eqs. (16) and (21) with $\Delta\Omega=\epsilon^{2}/2$, the timeperiod $T$ in the NLS approximation is
@@
where $\gamma(c)=\frac{\sqrt{-2c}}{6\sqrt{(2c+1)+\sqrt{8c+1}}}\frac{-5-4c-5\sqrt{8c+1}}{4c-1-\sqrt{8c+1}}$ and $\tilde{k}(c)\equiv\sqrt{\frac{-(2c+1)-\sqrt{8c+1}}{2c}}$ is determined by inverting $c=\partial_{\tilde{k}}\omega_{0}(\tilde{k},1)$. All NLS-like envelope solitons that enclose several carrier wave oscillations within the breather core, lie close to the two-parameter surface $T=\mathcal{T}(\tilde{a},c)$. These breather cores decay rapidly to a very small-amplitude ($a\ll 1$) periodic background. This surface is depicted in Fig. 6(b). The parameters $(\tilde{a},c,\;{\rm and}\;T)$ associated with each computed traveling breather are also rendered in the figure. In the weakly nonlinear regime, the traveling breathers of all five families reside close to $T=\mathcal{T}(\tilde{a},c)$. Eventually, they depart from the surface, exhibiting a larger time period than weakly nonlinear theory predicts. For families $1$-$4$, the strongly nonlinear breathers limit to enclosing very few carrier oscillations within the breather cores, with the envelope widths being comparable to the cnoidal (background) carrier wavelength. An exception in this regard is family 5; wherein the breather core possesses a slowly varying envelope despite a large amplitude $\tilde{a}$. As a final remark, we draw attention to families 2 and 3, computed on domains with $L=100$ and $250$ respectively, and seeded with an identical NLS envelope soliton (see Fig. 6). The role of the computational domain length as an additional parameter (besides $c$ and $T$) is clear, as the scatter plots diverge noticeably beyond the weakly nonlinear regime owing to the different induced phase shift $\sigma$.